two dice are rolled
A) what is the probability of rolling more than a 9?
B) what is the probability of rolling less than 4?
C) what is the probability of rolling 13?

Answer by Boreal(15235)

2-1
3-2
4-3
5-4
6-5
7-6
8-5
9-4
10-3
11-2
12-1
There are 36 outcomes
More than a 9 means a 10, 11, or a 12, or 6/36=(1/6)
Less than a 4 means a 3 or a 2 (3/36)=(1/12)
A 13 has probability 0. It cannot happen